### v2

# DexDiffuser: Generating Dexterous Grasps with Diffusion Models

Abstract
We introduce DexDiffuser, a novel dexterous grasping method that generates, evaluates, and refines grasps on partial object point clouds. DexDiffuser includes the conditional diffusion-based grasp sampler DexSampler and the dexterous grasp evaluator DexEvaluator. DexSampler generates high-quality grasps conditioned on object point clouds by iterative denoising of randomly sampled grasps. We also introduce two grasp refinement strategies: Evaluator-Guided Diffusion and Evaluator-based Sampling Refinement. The experiment results demonstrate that DexDiffuser consistently outperforms the state-of-the-art multi-finger grasp generation method FFHNet with an, on average, 9.12% and 19.44% higher grasp success rate in simulation and real robot experiments, respectively. Supplementary materials are available at https://yulihn.github.io/DexDiffuser_page/
I Introduction
Despite years of research on data-driven grasping [1], generating dexterous grasps for picking unknown objects, similar to the example shown in Fig. LABEL:fig:pull_figure, remains challenging [2]. One of the main challenges in dexterous grasping is identifying successful grasps within a high-dimensional search space [3].
In this work, we propose the new data-driven dexterous grasping method DexDiffuser to address the dimensionality issue. DexDiffuser includes DexSampler, a new conditional diffusion-based dexterous grasp sampler, and DexEvaluator, a new dexterous grasp evaluator. DexSampler generates high-quality grasps conditioned on object point clouds by iteratively denoising randomly sampled grasps. In addition to DexDiffuser, we also propose two grasp refinement strategies: Evaluator-Guided Diffusion (EGD), which is only applicable for diffusion models, and Evaluator-based Sampling Refinement (ESR), which applies to all dexterous grasp sampling methods.
We experimentally assess DexDiffuser’s capability to sample, evaluate, and refine 16- Degrees of Freedom (DoF) Allegro Hand grasps on three visually and geometrically distinctive object datasets in simulation and on one dataset in the real world. DexDiffuser is benchmarked against the state-of-the-art multi-finger grasp generation method FFHNet [2]. The experimental results indicate that our best method achieves a grasp success rate of 98.77% in simulation and 68.89% in the real world, which is 11.63% and 20.00% higher than the respectively best FFHNet model.
Our contribution can be summarized as follows:
  • DexSampler the novel diffusion-based dexterous grasp sampler (Section IV-B), and DexEvaluator the dexterous grasp evaluator (Section IV-C).
  • The two refinement strategies EGD and ESR for improving dexterous grasps (Section IV-D).
  • A dataset of 1.7 million successful and unsuccessful Allegro Hand grasp across 5378 objects (Section V).
  • A comprehensive experimental evaluation demonstrating the efficacy and real-world applicability of DexDiffuser (Section VI).
II Related Work
This work spans two topics: data-driven dexterous grasping and diffusion models. For brevity, we limit the related works on diffusion models to robotics and, for other areas, refer to the comprehensive survey by Yang et al. [4].
II-A Data-Driven Dexterous Grasping
Data-driven dexterous grasp sampling methods primarily fall into three categories: 1) those that generate a contact map on the object’s surface [5, 6, 7], 2) those that rely on shape completion [8, 9, 10, 11, 12, 13, 14, 15], or 3) those that train a grasping policy based on Reinforcement Learning (RL) [16, 17] or human demonstrations [18]. The main limitation of the methods in 1) is the assumption of complete object observation, which hinders real-world applications where complete visibility cannot be guaranteed. Methods in 2) are limited because the generated grasp quality is closely intertwined with the completion quality while collecting expensive demonstrations, and the sim-to-real gap limits the methods in 3).
Only a few methods [19, 20, 2] do not adhere to any of the above categories as they generate grasp poses directly on incomplete object observations without shape completion nor training a grasping policy. Our early work [19], focuses on building hierarchical representations of objects on which grasps are sampled. To make learning more tractable, Choi et al. [20] considered one gripper configuration and discretized grasp poses into 6 approach directions and 4 orientations. A discriminative network is trained to score the 24 pre-selected grasp poses for the highest grasp success probability. Qian et al. [2] proposed FFHNet: a Variational Autoencoder (VAE)-based grasp sampler that generates 15-DoF gripper configurations and grasp poses in all of SE(3). Albeit FFHNet reached a 91% grasp success rate in simulation, it was never evaluated on real hardware, leaving the question of real-world performance unanswered. DexDiffuser also produces complete gripper configurations and SE(3) grasp poses but by using diffusion models. Also, we experimentally validate DexDiffuser and FFHNet on real robotic hardware.
II-B Diffusion Models in Robotics
Despite beind rather recent, diffusion models have already found widespread applicability in motion planning [21], navigation [22], manipulation [23, 24, 25, 26], human-robot interaction [27] and grasping [28]. Of these, the most related to ours is the grasping work by Urain et al. [28], where a diffusion model was trained to generate SE(3) parallel-jaw grasp poses. We propose a diffusion model for generating dexterous grasps, meaning that both SE(3) grasp poses and high dimensional gripper configurations are generated.
Prior works on diffusion models in robotics have also trained discriminators to evaluate the diffusion-generated samples [24, 25]. In [24], the discriminator scores how realistic point cloud-generated scenes are, while the discriminator in [25] scores how good a generated SE(3) pose is for manipulating an object. Similarly, we use a discriminator to evaluate a generated grasp’s success probability in picking an object.
III Problem Statement
We consider the problem of sampling and evaluating dexterous grasps $\mathbf{G}$ on partially observed point clouds $\mathbf{O}\in\mathbb{R}^{\text{N}\times 3}$, where $N$ is the number of points, for the task of successfully picking up objects. In this work, a grasp $\mathbf{g}\in\mathbf{G}=[\mathbf{p},\mathbf{r},\mathbf{q}]\in\mathbb{R}^{9+k}$ is represented by a 3-D position $\mathbf{p}\in\mathbb{R}^{3}$, a continouous 6-D rotation vector $\mathbf{r}\in\mathbb{R}^{6}$ [29], and a gripper joint configuration $\mathbf{q}\in\mathbb{R}^{k}$, where $k$ is the number of controllable joints. Henceforth, $\mathbf{p}$ and $\mathbf{r}$ are together referred to as the grasp pose and $\mathbf{q}$ as the joint angles. We assume all objects are reachable and singulated, and the object point cloud is segmented from the scene.
We frame the problem of sampling grasps as learning a distribution $\mathsf{P}{(\mathbf{g}\in\mathbf{G}|\mathbf{O})}$ and of evaluating grasps as learning a discriminator $\mathsf{P}(S=1|\mathbf{g},\mathbf{O})$, where $S$ is a binary variable representing grasp success ($S$=1) and failure ($S$=0). Henceforth, $\mathsf{P}{(S=1|\mathbf{g}\in\mathbf{G},\mathbf{O})}$ is referred to as the grasp evaluator and $\mathsf{P}{(\mathbf{g}\in\mathbf{G}|\mathbf{O})}$ as the grasp generator. We approximate each of these distributions with a separate Deep Neural Network (DNN) $\mathcal{D}_{\bm{\theta}}(\mathbf{g}\in\mathbf{G}|\mathbf{O})\approx\mathsf{P}{(\mathbf{g}\in\mathbf{G}|\mathbf{O})}$ and $\mathcal{D}_{\bm{\psi}}(S=1|\mathbf{g}\in\mathbf{G},\mathbf{O})\approx\mathsf{P}{(S=1|\mathbf{g}\in\mathbf{G},\mathbf{O})}$, with learnable parameters $\bm{\theta}$ and $\bm{\psi}$. The objective then becomes learning $\bm{\theta}$ and $\bm{\psi}$ from data.
IV Method
We now introduce our framework DexDiffuser to learn the parameters $\bm{\theta}$ and $\bm{\psi}$. DexDiffuser consists of two parts: DexSampler and DexEvaluator. We also propose two grasp refinement strategies for improving the grasp success probability.
IV-A The Basis Point Set Representation
DexDiffuser generates and samples grasps directly on partial object point clouds. In this work, $\mathbf{O}$ is encoded into $\mathbf{f}_{\mathbf{O}}$ using Basis Point Set (BPS) [30] as this encoding has shown to work well in prior dexterous grasping work [2]. The BPS encoding represents the shortest distance between each point in a fixed set of basis points and all points in $\mathbf{O}$. The benefit of BPS is that it captures the shape and spatial properties of the object and is of fixed length, which allows for the use of computationally efficient Fully Connected (FC) layers to further process $\mathbf{f_{O}}$.
IV-B Diffusion-based Grasp Sampler
Figure 1: Model architecture. The left image shows the DexSampler. Its input $\mathbf{f_{O}}$ is processed into a key-value pair while $\mathbf{g}_{t}$ and $t$ are processed into a query using a self-attention block. The key-value-query triplet is then embedded using a cross-attention block to compute $\hat{\mathbf{\epsilon}}_{t}$. The right image shows the DexEvaluator that predicts grasp success probability given the same $\mathbf{f_{O}}$ as the DexSampler and $\mathbf{g}_{0}$ produced by the DexSampler.
DexSampler is a classifier-free, conditional diffusion model that generates $\mathbf{G}$ conditioned on $\mathbf{f_{O}}$. It consists of a forward process that gradually turns data into Gaussian noise and a learnable inverse diffusion process that recovers the data from noise given the condition [31].
In our case, the forward process for incrementally diffusing a successful grasp $\mathbf{g}_{0}=[\mathbf{r}_{0},\mathbf{p}_{0},\mathbf{q}_{0}]$ over $T$ timesteps is
$q(\mathbf{g}_{1:T}|\mathbf{g}_{0})$ $=\Pi_{t=1}^{T}q(\mathbf{g}{{}_{t}}|\mathbf{g}_{t-1}),$ (1)
$q(\mathbf{g}{{}_{t}}|\mathbf{g}{{}_{t-1}})$ $=\mathcal{N}(\mathbf{g}{{}_{t}};\sqrt{1-\beta_{t}}\mathbf{g}{{}_{t-1}},\beta_{t}\mathbf{I}),$ (2)
where $\beta_{t}$ is the scheduled noise variance at time step $t$. To recover the original $\mathbf{g}_{0}$ from $\mathbf{g}_{T}$, DexSampler learns the reverse diffusion steps by estimating the Gaussian noise added at each $t$ and then removes it as
$p_{\bm{\theta}}(\mathbf{g}{{}_{0:T}}|\mathbf{f_{O}})$ $=p(\mathbf{g}{{}_{T}})\Pi_{t=1}^{T}p_{\bm{\theta}}(\mathbf{g}{{}_{t-1}}|\mathbf{g}{{}_{t}},\mathbf{f_{O}}),$ (3)
$p_{\bm{\theta}}(\mathbf{g}{{}_{t-1}}|\mathbf{g}{{}_{t}},\mathbf{f_{O}})$ $=\mathcal{N}(\mathbf{g}{{}_{t-1}};\hat{\mu}_{\bm{\theta}}(\mathbf{g}_{t},t,\mathbf{f}{{}_{O}}),\hat{\Sigma}_{{\bm{\theta}}}(\mathbf{g}_{t},t,\mathbf{f}{{}_{O}})),$ (4)
where $\hat{\mu}_{\bm{\theta}}(\mathbf{g}_{t},t,\mathbf{f}{{}_{O}})$ and $\hat{\Sigma}_{{\bm{\theta}}}(\mathbf{g}_{t},t,\mathbf{f}{{}_{O}})$ is the learnable mean and variance of a Gaussian distribution. The parameters $\bm{\theta}$ are learned by minimizing the loss function
$L_{\epsilon}=||\hat{\mathbf{\epsilon}}{{}_{t}}-\mathbf{\epsilon}{{}_{t}}||^{2},$ (5)
where $\hat{\mathbf{\epsilon}}{{}_{t}}\sim\mathcal{N}(\mathbf{g}{{}_{t-1}};\hat{\mu}_{\bm{\theta}}(\mathbf{g}_{t},t,\mathbf{f}{{}_{O}}),\hat{\Sigma}_{{\bm{\theta}}}(\mathbf{g}_{t},t,\mathbf{f}{{}_{O}}))$ and $\mathbf{\epsilon}_{t}$ is the ground-truth noise.
The architecture of DexSampler is depicted in Fig. 1, following the practical design of conditional diffusion model [32, 33, 34, 35]. It takes three inputs: $\mathbf{f_{O}}$, $\mathbf{g}_{t}$, and $t$ and outputs $\mathbf{\hat{\epsilon}}_{t}$ for grasp denoising.
IV-C Grasp Evaluator
Although DexDiffuser is trained to generate successful grasps, it can still produce unsuccessful ones. Thus, it would be useful if we could only execute grasps with a high probability of succeeding. To this end, we propose DexEvaluator $\mathcal{D}_{\bm{\psi}}(S=1|\mathbf{g}\in\mathbf{G},\mathbf{f_{O}})$ to predict the grasp success probability. The architecture of DexEvaluator is shown in Fig. 1 and is inspired by the evaluator in [2]. It takes two inputs: $\mathbf{g}$ and $\mathbf{f_{O}}$ and outputs $S$. The main difference to [2] is that we introduce frequency encoding [36] that maps $\mathbf{g}$ into a higher dimensional space before passing it to the evaluator. The map takes the following form
$\varphi(\mathbf{x})=$ $[\sin(2^{0}\pi\mathbf{x}),~{}\cos(2^{0}\pi\mathbf{x}),~{}\ldots,$
$\sin(2^{\text{F}-1}\pi\mathbf{x}),~{}\cos(2^{\text{F}-1}\pi\mathbf{x})],$ (6)
where F specifies the number of sinusoidal functions. $\varphi(\cdot)$ is then separately applied on $\mathbf{p}$, $\mathbf{r}$, and $\mathbf{q}$. Introducing frequency mapping into the evaluator allows it to model better how small changes in $\mathbf{g}$ can lead to large changes in $S$ [37], which requires the model to be sensitive to high-frequency details [36].
The parameters $\bm{\psi}$ of the DexEvaluator is optimized using the binary cross-entropy loss
$L(y,\hat{y})=-y\log(\hat{y})+(1-y)\log(1-\hat{y}),$ (7)
where $y$ and $\hat{y}$ denote the ground truth and predicted grasp labels, respectively. In Section V, we provide details on the dataset used to train DexEvaluator.
In addition to evaluating grasps, the DexEvaluator can also be used to guide the inverse diffusion process towards more successful grasps or refine already sampled grasps. This we discuss next.
IV-D Grasp Refinement
The idea behind grasp refinement is to increase the probability of grasp success. In this work, we propose two grasp refinement strategies: EGD, which uses the evaluator to guide the inverse diffusion process, and ESR, which uses the evaluator to refine already sampled grasps.
IV-D1 EGD
EGD is a form of classifier-guided diffusion. Formally, during the inverse diffusion process, a noisy grasp $\mathbf{g}{{}_{t}}$ is iteratively denoised as
$p_{\theta}(\mathbf{g}{{}_{t-1}}|\mathbf{g}{{}_{t}},\mathbf{f_{O}})=$ $\mathcal{N}\left(\mathbf{g}{{}_{t-1}};\mu_{\theta}(\mathbf{g}{{}_{t}},t,\mathbf{f}{{}_{O}})+\lambda\cdot\mathbf{g}{}{{}_{\text{guide}}},\right.$
$\Sigma_{\theta}(\mathbf{g}{{}_{t}},t,\mathbf{f_{O}})\right),$ (8)
$\mathbf{g}{{}_{\text{guide}}}=$ $\nabla_{\mathbf{g}{{}_{t}}}\log\mathcal{D}_{\phi}(\hat{y}=1|\mathbf{g}{{}_{t}},\mathbf{f_{O}})+\log(T-t+1),$ (9)
where $\mathbf{g}{{}_{\text{guide}}}$ is the gradient signal from DexEvaluator modulated by $\log(T-t+1)$. In each step of the inverse diffusion process, $\mathbf{g}{{}_{\text{guide}}}$ is used to adjust $\mu_{\theta}$ towards more successful grasp. The magnitude of the adjustment is controlled by the parameter $\lambda$. The time-dependent modulation term $\log(T-t+1)$ ensures that the guidance is more pronounced in the earlier stages of denoising when the uncertainty in the grasp is higher.
IV-D2 ESR
Compared to EGD, ESR locally refines already sampled grasps. Prior research on parallel-jaw grasp sampling [37] has demonstrated that many sampled grasps with low evaluator scores often lie spatially close to higher-scoring ones, suggesting that local grasp refinement can improve the grasp success probability. Building upon this insight, we extend the concept of local refinement for parallel-jaw grasps [37, 38] to dexterous grasps, which requires us to refine $\mathbf{p}$, $\mathbf{r}$ and $\mathbf{q}$.
Mathematically, local grasp refinement boils down to finding a $\Delta\mathbf{g}$ so that $\mathsf{P}{(S=1|\mathbf{g}+\Delta\mathbf{g},~{}\mathbf{f_{O}})}>\mathsf{P}{(S=1|\mathbf{g},~{}\mathbf{f_{O}})}$. To find $\Delta\mathbf{g}$, we use the Metropolis-Hasting (MH) sampling algorithm as it has been shown to perform well for refining parallel-jaw grasps [38]. To apply MH, we first sample $\Delta\mathbf{g}$ from a known probability distribution and then calculate the acceptance ratio $\alpha=\frac{\mathcal{D}_{\bm{\psi}}(S=1|\mathbf{g}+\Delta\mathbf{g},\mathbf{f_{O}})}{\mathcal{D}_{\bm{\psi}}(S=1|\mathbf{g},\mathbf{f_{O}})}$. If $\alpha\geq u$, where $u\sim\mathcal{U}[0,1]$, then $\mathbf{g}=\mathbf{g}+\Delta\mathbf{g}$ otherwise $\mathbf{g}$ is not updated. We refer to this process as ESR-1, as it performs refinement simultaneously for every element of $\mathbf{g}$.
Because of the higher dimensional search space, local refinement is more challenging for dexterous than parallel-jaw grasps. Therefore, to make the refinement process more tractable, we introduce a two-stage ESR (ESR-2) that first refines $\mathbf{p}$ and $\mathbf{r}$ jointly and only afterward refines $\mathbf{q}$. The idea behind ESR-2 is first to refine the global grasp parameters $\mathbf{p}$ and $\mathbf{r}$ and only when these are good, refine the local parameters $\mathbf{q}$.
IV-E Implementation Details
[…]
For calculating the BPS we randomly sample one set of 4096 basis points. This set is then used to calculate the BPS for every $\mathbf{O}$ in the dataset.
V Dataset
We used DexGraspNet [39] to generate training grasps and Isaacgym to render 50 $\mathbf{O}$ from different camera views for each object. We collected over 0.7M successful and 0.8M unsuccessful Allegro Hand grasps on 5378 objects of various geometries and scales. Only successful grasps were used to train DexSampler while both successful and unsuccessful grasps were used to train DexEvaluator.
To enhance the dataset with out-of-distribution negative samples, we generated more grasp candidates by jittering the successful grasps in the original dataset. Specifically, a grasp perturbation $\Delta\mathbf{g}_{k}^{\text{{pert}}}$ was added for every successful object-grasp pair $(\mathbf{O}_{k},\mathbf{g}_{k})$. We tested and labeled all the perturbed grasps in Isaacgym and only kept the unsuccessful ones.
The final dataset consisted of 1.7 million grasps, where approximately 40.52% were successful and 59.48% were unsuccessful. We split this dataset into a training set of 4303 objects and a test set of 1,075. We further add 49 objects from the EGAD! dataset [40] and 16 objects from the MultiDex dataset [5] to the the test dataset.
VI Experimental evaluation
We experimentally evaluated DexDiffuser in both simulation and the real world, comparing it to the state-of-the-art grasp sampler and evaluator FFHNet [2]. The specific questions we wanted to address with the experiments were:
  1. 1.
    How well can the DexEvaluator predict grasp success?
  2. 2.
    What effect does the BPS-encoding have on generating successful grasps?
  3. 3.
    To what extent do ESR and EGD improve the grasp quality?
  4. 4.
    How well can DexDiffuser generalize to real-world object grasping?
VI-A Simulation Experiments
All simulation experiments were conducted in Isaac Gym [41] following the setup from [5]. Specifically, test objects were initialized as free-floating objects, and a virtual 16-DoF Allegro Hand was used for grasping each of them. For each of the 1140 test objects, 10 point clouds were rendered from randomly sampled camera poses. Then, all methods generated 20 grasps per point cloud, which amounted to 228,000 test grasps in total. Once an object was grasped, the grasp stability was assessed by applying external forces along the six directions in space for 50 timesteps. A grasp succeeded if the object remained in the hand throughout the stability test.
We also trained a DexSampler with the PointNet++ point cloud encoding [42] to study how different encodings affect grasp performance. We call that network DexDiffuser-PN2 while the BPS network is referred to as DexDiffuser-BPS.
Figure 2: Generated grasp examples. Qualitative evaluation of DexSampler-BPS on objects from different datasets. The grasps, shown in pink, are generated on the partially observed point clouds, shown in red, rendered from the complete meshes, shown in blue.
{adjustbox}
max width= Evaluator Recall Pos. (%$\uparrow$ Recall Neg. (%$\uparrow$ Total Acc. (%$\uparrow$ No Freq. Enc. 56.47 90.88 73.68 $(10,0,0)$ 77.28 83.42 80.35 $(0,10,0)$ 54.19 91.35 72.77 $(0,0,10)$ 77.13 78.22 77.65 $(5,5,5)$ 55.49 91.78 73.64 $(10,10,10)$ 75.92 78.01 76.97 $(10,4,4)$ 55.32 91.54 73.43 $(20,4,0)$ 74.53 83.68 79.10 $(10,4,0)$ 80.85 80.97 80.91
TABLE I: Grasp success prediction. $(f_{1},f_{2},f_{3})$ represent $F$ in (IV-C) for $\mathbf{p}$, $\mathbf{r}$, and $\mathbf{q}$, respectively. $\uparrow$: The higher, the better.
{adjustbox}
max width= Method Dataset MultiDex(%$\uparrow$ EGAD! (%$\uparrow$ DexGN (%$\uparrow$ FFHNet [2] 72.47 87.14 82.16 DexS-PN2 82.28 98.16 81.77 DexS-BPS 87.28 98.62 82.64 FFHNet-ESR-1 71.91 86.77 82.05 FFHNet-ESR-2 72.16 86.86 82.21 DexS-BPS-ESR-1 87.16 98.67 82.57 DexS-BPS-ESR-2 87.47 98.68 82.60 DexS-BPS-EGD 87.63 98.68 82.88 DexS-BPS-EGD-ESR-1 87.41 98.68 82.94 DexS-BPS-EGD-ESR-2 87.78 98.77 82.90
TABLE II: Grasp success rate of different methods. The three first rows exclude grasp refinement, while the bottom nine include it. DexS represents the DexSampler. DexGN is short for the DexGraspNet test split. $\uparrow$: The higher, the better.
{adjustbox}
max width= Real-world experiment # of successful grasps per method and object Object FFHNet FFHNet-ESR-2 DexS-BPS DexS-BPS-ESR-2 DexS-BPS-EGD DexS-BPS-EGD-ESR-2 1. Crackerbox 2/5 4/5 4/5 5/5 4/5 4/5 2. Sugar box 3/5 4/5 4/5 4/5 4/5 5/5 3. Mustard bottle 3/5 3/5 4/5 5/5 3/5 3/5 4. Bleach cleanser 2/5 2/5 4/5 4/5 5/5 4/5 5. Sprayer 2/5 2/5 3/5 4/5 4/5 4/5 6. Metal mug 2/5 1/5 2/5 3/5 1/5 3/5 7. Goblet 3/5 3/5 5/5 5/5 5/5 4/5 8. Toy plane 0/5 0/5 0/5 0/5 0/5 0/5 9. Pringles 2/5 3/5 2/5 1/5 3/5 2/5 Avg. success rate 42.22% 48.89% 62.22% 68.89% 64.44% 64.44% Time (ms) 2 490 1692 2188 2548 3036
TABLE III: Experimental results for real-world grasping. DexS represents DexSampler.
VI-A1 Grasp Evaluation
To answer the first question, we assessed the DexEvaluator’s grasp prediction accuracy over different numbers of position encodings. The results are presented in Table I. The findings indicate that the DexEvaluator configured with frequencies $(10,4,0)$ achieves the best performance and maintains a balanced accuracy between positive and negative predictions. DexEvaluator with few or no frequencies is biased towards negative predictions, while increasing the number of frequencies from 10 to 20 reduces performance. Thus, we choose the $(10,4,0)$ DexEvaluator in all subsequent experiments.
VI-A2 Grasp Sampling
To address the second question, we evaluated both FFHNet and DexSampler using either the PointNet++ or BPS encoding. The simulation results are presented in Table II. These results demonstrate that DexSampler, irrespective of the encoding method, consistently outperformed the FFHNet generator across all test datasets. The superior performance underscores the diffusion models’ effectiveness in synthesizing successful grasps. Moreover, DexSampler-BPS performed much better than DexSampler-PN2, indicating that the BPS encoding is less affected by point cloud irregularities, including missing parts and non-uniform density distribution. Finally, Fig. 2 shows some grasps generated by the DexSampler-BPS on the three datasets.
VI-A3 Grasp Refinement
To answer question 3, we assessed how grasp refinement affected the grasp success rate. The results, presented in Table II, show that ESR improves grasp success rate for both DexDiffuser and FFHNet. For the diffusion-based models, EGD showed marginal improvements, consistent with findings in diffusion-based robot rearrangement scenarios [43]. The two-stage refinement process ESR-2 led to slightly better grasp success rates than ESR-1, giving some evidence to support our hypothesis that it is better to refine global grasp parameters before local ones.
We also evaluated the combination of EGD and ESR (EGD-ESR) on DexDiffuser by first applying EGD followed by ESR. Although the combination did achieve the highest grasp success rate on the MultiDex and EGAD! datasets, the improvement is, again, marginal compared to the other methods. We hypothesize the reason refinement does not increase the grasp success rates more is because the average grasp classification accuracy, even for the best model, is only 80.91%.
VI-B Real Robot Experiment
To answer the last question, we evaluated how successful DexDiffuser was at picking objects from real noisy object point clouds using real robotic hardware. The hardware used is shown in Fig. LABEL:fig:pull_figure and includes an Allegro Hand attached to the end of a Franka Panda robot and a Kinect V3 placed in a third-person perspective to capture $\mathbf{O}$. We used a motion capture system to do the extrinsic camera calibration.
We chose 9 different test objects, all shown in Fig. 3, based on their variation in shape and size. Each object was placed in the robot’s workspace in five distinct orientations: 0°, 72°, 144°, 216°, and 288°. All methods generated and ranked 200 grasps per object and orientation. Of these grasps, the best kinematically feasible one was executed on the robot. A grasp was successful if the robot picked the object and kept it grasped while moving to a predefined pose where it rotated the last joint $\pm$90°.
Figure 3: Experimental objects. From 1 to 9: cracker box, sugar box, mustard bottle, bleach cleanser, sprayer, metal mug, goblet, toy plane, and pringles
Figure 4: Successful and failed grasp examples from DexSampler-BPS-ESR-2. The first two images successfully grasp the bleach cleanser and the goblet, while the latter two show different failure cases on the metal mug and the pringles.
We evaluated the following methods: FFHNet, FFHNet-ESR-2, DexDiffuser-BPS, DexDiffuser-BPS-ESR-2, DexDiffuser-BPS-EGD and DexDiffuser-BPS-EGD-ESR-2. It is worth highlighting that these methods are only trained on simulated data. The experimental results are summarized in Table III. These results, like the simulation results, show that DexDiffuser consistently outperformed FFHNet across all test objects. The best method for picking real-world objects was DexSampler-BPS-ESR-2, and some of the grasps it produced are shown in Fig. 4. Interestingly, no methods successfully grasped the toy plane, mainly because the Allegro hand was not strong enough to keep the object secure when the last joint rotated $\pm$90°, which is demonstrated in the accompanying video [^1] https://yulihn.github.io/DexDiffuser_page/.
The average grasp success rates in the real world are also consistently lower than in simulation. One rather apparent reason for the decline in performance is that in the real world, grasps are generated on noisy point clouds compared to the clean simulated ones. Another reason is the presence of obstacles that impede the robot’s ability to attain the desired target position. For instance, the first failed grasp in Fig. 4 is because the gripper collided with the table when closing, while the second failure is because the robot could not reach the intended grasp position due to joint limits. Neither of these failure cases is present in the simulation.
We also report the inference time in Table III. The iterative denoising process makes the diffusion-based methods slower at inference than FFHNet. For instance, the best performing DexDiffuser (DexS-BPS-ESR-2) is 3.5 times slower than the best performing FFHNet (FFHNet-ESR-2).
VII Conclusion
This paper addressed the problem of dexterous grasping under partial point cloud observations. Our proposed solution, DexDiffuser, consists of a conditional diffusion-based grasp sampler and grasp evaluator, BPS-encoded point clouds, and two different grasp refinement strategies. All experiments on the 16-DoF Allegro hand demonstrated that DexDiffuser outperformed FFHNet with an average of 9.12% and 19.44% higher grasp success rate in simulation and real robot experiments, respectively.
The limitations of DexDiffuser are the sim-to-real gap, not accounting for environmental constraints such as collision objects, and the relatively long grasp sampling time. Still, DexDiffuser is one of few data-driven dexterous grasping methods that generate grasps directly on partial point clouds and is evaluated on real hardware. Based on the improvements brought by the diffusion model, we believe such methods can be used to solve other challenging dexterous manipulation tasks, including in-hand manipulation, dexterous grasping in clutter, and joint motion and dexterous grasp planning. Addressing these challenges and the limitations of our method highlights promising future research directions.